# A Survey on Speech Large Language Models for Understanding

### v5

@@ IV-B1 Modality Fusion of Continuous Audio Embedding
    Pre-Decoder Alignment: This category refers to methods that align audio features with text embeddings before feeding them into the LLM decoder. A common approach is to project the audio feature information into the LLM’s text feature space through a connector [30, 26]. Specifically, the tensor containing audio features typically has its own distinct feature dimension. The audio feature embedding $Z_{a}$ is projected into a vector with the same dimensionality as the text embedding using a projection function $w$: $H_{a}=w(Z_{a})$, where $H_{a}$ denotes the projected audio embedding in the text embedding space. These audio embeddings are then concatenated with the input text embedding $H_{t}$ to form a new embedding sequence that integrates both speech and text information: $H=[H_{a},H_{t}]$. This combined sequence is then fed into the LLM. Some researchers have also implicitly incorporated the projection step within the original encoder, achieving modality projection by adjusting the encoder’s parameters during training [64]. Most models use a multi-layer perceptron (MLP) as the projection module, while some adopt more advanced techniques such as Q-Former, which introduces trainable query tokens that attend to audio features and output fixed-length embeddings aligned with the LLM input space [26].
@@ IV-B2 Modality Fusion of Discrete Audio Tokens
    Token Mapping: The most intuitive approach is to map audio tokens to text tokens that the LLM can process. Tsunoo et al. [86] propose using CTC predictions to efficiently generate prompts—referred to as CTC prompts. Once text tokens are obtained, they are combined with tokenized input text to form a unified token sequence containing both audio and text information. This sequence is then fed into the LLM for processing. This approach not only preserves the semantic content of the audio features but also ensures consistency in the LLM’s processing pipeline.
@@ IV-C LLM Inference
The LLM inference stage consists mainly of a text decoder, which is typically an autoregressive transformer decoder that generates text tokens conditioned on input audio and/or text. Architecturally, it often mirrors standard language models like GPT, but it is specially designed or adapted to handle multimodal inputs—such as projected audio embeddings or cross-attended audio features—alongside text tokens. To accommodate modality fusion, the decoder may incorporate special tokens (e.g., <aud>, <text>) or use prompt formatting to distinguish and contextualize different modalities during generation. Despite these adaptations, the decoder follows the same autoregressive objective: predicting the next text token given previous tokens and any additional input context. In ST2ST style models, the decoder may also be extended to support the generation of discrete audio tokens. One common approach, as seen in VALL-E, treats speech generation as a conditional codec language modeling task, where the model generates audio tokens conditioned on phoneme sequences and an acoustic prompt  [88, 57]. Another approach used in SpeechGPT expands the LLM’s vocabulary to include both text and speech tokens, enabling the model to generate audio token sequences directly alongside or following text, depending on the instruction[58]. In both cases, the generated audio tokens are passed through a neural codec decoder (e.g., EnCodec[55] or HiFi-GAN[89]) to reconstruct the final waveform.
@@ V Training Strategies
TABLE II: The following table shows the training strategy of Speech LLM, we show the training parameters at different stages, corresponding to the training data task type (as indicated inside the parenthesis) and the type of capabilities the final model has. The – symbol indicates that the paper does not mention it in detail, while ✘ symbol indicates that the model does not have this stage.
@@ V-A Modality Alignment
At the outset, Speech LLMs must learn a shared representation between spoken input and text. In this modality alignment phase, models typically leverage vast speech data to bridge the audio-text gap. A common strategy is to incorporate a pre-trained speech encoder and align its acoustic representations with a text-based LLM. For example, Qwen-Audio begins with a pretrained Whisper encoder coupled to a LLM backbone via a projector [92]. The alignment is refined by supervised tasks that force audio and text to correspond: notably, automatic speech recognition (ASR) and audio captioning (AAC) are usually used as fundamental alignment tasks [29, 27].
@@
The total parameters $\theta$ of the model are composed of $\theta_{\text{LLM}}$, $\theta_{\text{encoder}}$, and $\theta_{\text{adapter}}$ The model is trained to map an input audio sequence $a$ to a textual output $y$ (transcript or caption). The training data consist of many triples $(a,p,y)$: an audio input $a$, a textual prompt or task descriptor $p$, and the expected textual output $y$. The model is optimized to predict $y$ In this stage, the textual prompt $p$ typically remains unchanged, which makes modal alignment easier to train [66, 93]. The training loss can be represented as
$\mathcal{L}_{\text{align}}(\theta)=-\log P(y\mid\theta_{\text{LLM}}(\theta_{\text{adatper}}(\theta_{\text{encoder}}(a)),p).$ (3)
@@
Depending on the training objectives, many models focus on single speech-to-text task (e.g., speech recognition) rather than training a multitask speech model capable of fully understanding speech. Therefore, many models only have one stage of modality alignment. The most common approach is to train only the adapter [94, 93]. However, after the initial alignment, continuing to add LoRA parameters for further training has proven to maintain better modality alignment and achieve superior alignment performance [66]. Another training approach involves simultaneously training the encoder, adapter, and LoRA parameters, as seen in SLAM and FireRedASR [95, 31]. Interestingly, both models employ Conformer as the encoder, and the FireRedASR model has achieved SOTA performance on current recognition leaderboards—perhaps demonstrating that unfreezing the encoder and training with more parameters can yield better alignment results. There is currently no unified consensus on the order of training parameters. A common practice is to first train the adapter, then train the encoder and adapter, and finally add LoRA parameters for joint training [33]. However, it is also advisable to include stages where some parameters are frozen during training [96]. For multitask models speech like Qwen-audio [64] and Seed-ASR [27], LoRA parameters of LLMs are typically not used as training parameters for modality alignment—partly to preserve the original powerful contextual and reasoning abilities of the LLMs, expecting them to play a role in subsequent multitask training. Not all models follow this pattern, however: SALMONN [26] trains LoRA parameters during the alignment phase.
@@ V-B Multitask Training
After modality alignment, speech features are mapped into the text space of the LLM. However, even when a LLM can accurately transcribe speech, it often loses the ability to reason about or follow instructions related to the transcribed content. Therefore, the LLM must undergo a multitask training phase to become a multitask speech model.
Taking models focused solely on ASR tasks as an example, during this phase, many models add LoRA parameters to leverage the LLM’s contextual capabilities—such as using hotwords or historical transcripts—to improve recognition results [97, 27]. However, these are not standard multitask training approaches but rather extensions of ASR tasks.
@@
The mainstream training approach resembles modality alignment: natural language instructions are concatenated with the aligned speech features and target text and the model is directly trained for multitask via cross-entropy loss . Generally, LoRA parameters are added to the training at this stage to fine-tune the LLM, aiming to enable the LLM to acquire the multitask speech capability .
Notably, Qwen [64] employs a Whisper-like multitask training framework that uses task prefixes and does not use LoRA for fine-tuning. In contrast, OSUM [33] uses an ASR+X approach, where X represents other tasks. This means OSUM transcribes speech into text before performing any other audio tasks.
@@ V-C Preference Alignment
After multitask alignment, some advanced Speech LLMs apply preference alignment using reinforcement learning  [92]. This step optimizes the model’s responses according to human feedback or a learned reward function, refining qualities such as factual accuracy, relevance, or user satisfaction [98, 99], for example Seed-ASR[27]. A prominent method is reinforcement learning from human feedback (RLHF) operationalized via Proximal Policy Optimization (PPO) [100]. In RLHF, the model is treated as a policy $\pi_{\theta}(y|X)$ that generates a response $y$ to an input $X$, and a reward model $R(X,y)$ (trained from human preference rankings) scores the response. The goal is to maximize expected reward $E_{X}[E_{y\sim\pi_{\theta}}[R(X,y)]]$ while keeping the language model’s output distribution close to the pre-trained model to avoid degrading fluency. PPO achieves this by updating $\theta$ in the direction of the policy gradient $\nabla_{\theta}J\approx E_{X,y}[R(X,y)\,\nabla_{\theta}\log\pi_{\theta}(y|X)]$ with constraints on the KL-divergence from the original policy. In practice, the PPO update uses a clipped surrogate objective to ensure stable improvements. An alternative, lightweight approach is Direct Preference Optimization (DPO) [101], which was employed in Qwen2-Audio’s final training. Unlike methods that require training a separate reward model, DPO directly leverages preference pairs $(y^{+},y^{-})$ for a given prompt $x$. It optimizes the policy to assign higher probability to the preferred response $y^{+}$ over the less favored $y^{-}$. The DPO loss can be formulated as:
@@
where $\sigma$ is the sigmoid function and $\beta$ is a temperature hyperparameter. By minimizing this loss, the model is encouraged to prefer $y^{+}$ over $y^{-}$, effectively aligning its outputs with human preferences without the complexity of a reward model training loop. By the end of this reinforcement or preference alignment phase, the Speech LLM is highly tuned to follow instructions in a human-preferred way: it will listen to a user’s voice request and respond with appropriate content. Crucially, this stage often improves the model’s factuality and safety (by penalizing incorrect or undesirable continuations) – Qwen2-Audio’s authors note that after DPO fine-tuning, the model more faithfully adheres to user intent and exhibits fewer mistakes in complex audio-centric instructions. The strength of the instruction alignment phase lies in making the model not just capable, but also reliably usable: it transforms the multitask-trained foundation into a polite conversational agent that can handle open-ended voice interactions. However, it is worth noting that this phase has received limited attention in current Speech LLM research focused on speech understanding. Most existing works primarily emphasize modality alignment and task-specific performance, while instruction-following behavior, dialogue consistency, and safety alignment remain underexplored. As Speech LLMs continue to evolve toward general-purpose audio agents, further investigation into instruction and preference alignment is expected to play a crucial role in improving usability, controllability, and user interaction quality.
@@ V-D Training of Speech LLMs with Discrete Sequence Modeling
In contrast to the three-stage training pipeline of continuous embedding models, Speech LLMs based on discrete speech tokens adopt a simpler and more unified training strategy. For these models (e.g., AudioPaLM [59], SpeechGPT [58], VALL-E [88]), after generating the discrete tokens for audio representations, these token sequences are then treated analogously to text tokens, allowing a pretrained or newly initialized language model to be trained directly using a standard autoregressive objective.
Training typically proceeds by fine-tuning on paired speech-text data in a unified token space, without the need for a separate modality alignment stage, since discrete-token models do not introduce structural modifications such as adapters, which are common in continuous embedding models. This simplifies the pipeline and enables models to benefit directly from pretrained LLMs. In practice, alignment between modalities is learned implicitly through multitask fine-tuning, where the model learns to interleave and reason over both speech and text tokens. For example, SpeechGPT[58] introduces a “Paired Speech-Text Pretraining” phase to warm up the model on interleaved sequences, followed by instruction fine-tuning using synthetic multimodal dialogues. Unlike continuous-embedding Speech LLMs, mainstream discrete-token Speech LLMs typically skip explicit preference alignment (e.g., RLHF or DPO). Their training focuses on supervised multitask fine-tuning over speech–text data without a separate reward modeling stage. While emerging research such as SpeechAlign[102] has begun to explore preference optimization for codec-based language models, these methods remain experimental and are not yet part of the standard speech‑LLM training toolbox.
@@ VI Datasests
In parallel with model advancements, the availability and design of datasets have also evolved to support increasingly complex speech understanding tasks. In this section, we categorize and analyze these datasets through the lens of their respective understanding objectives, as defined in II-B. We categorize speech understanding tasks based on their objectives, ranging from simpler, more immediate tasks—such as perception tasks—to more complex and deeper objectives that necessitate advanced reasoning, including shallow and deep cognition tasks. This progression parallels the evolution of speech models, which initially focused on handling fundamental tasks but have since advanced to address more intricate cognitive functions. Here, we provide a detailed overview of the various categories of speech datasets and examine how they have evolved, particularly in response to the development of LLMs and TTS.
@@ VI-A Datasets for Perception and Shallow Cognition Tasks
Perception and shallow cognition tasks focus on extracting information from audio signals and performing intuitive understanding, without requiring deep semantic reasoning. Tasks such as Automatic Speech Recognition (ASR), speaker diarization, and keyword spotting are examples of perception tasks, while emotion recognition and speech translation are considered shallow cognition tasks. These types of tasks correspond to traditional single-task datasets that predate the era of LLMs. A significant portion of the datasets used to train Speech LLMs consists of audio-text pairs, which are readily available from many traditional single-task datasets. In addition to these audio-text pairs, tasks like speaker diarization and keyword spotting often include additional metadata, such as timestamps and speaker labels, to enhance task performance. Rather than being limited to single-task datasets, multi-task datasets with rich metadata are also widely used, enabling training across multiple tasks. The AMI Meeting Corpus is a representative example of such datasets, which can be used for a variety of tasks as shown in [128, 129, 130, 131].These datasets have become foundational in speech understanding due to their early development and wide availability. In addition to being used for pre-training Speech LLMs, these datasets can also serve as direct resources for downstream tasks, facilitating the fine-tuning of Speech LLMs.
Table III categorizes datasets associated with perception and shallow cognition tasks, including ASR, speaker analysis, keyword spotting, emotion classification, and speech translation. This categorization highlights the diversity of task types, language coverage, and dataset scale, reflecting the fundamental role of these resources in the broader field of speech understanding.

### v6

@@ IV-B1 Modality Fusion of Continuous Audio Embedding
    Pre-Decoder Alignment: This category refers to methods that align audio features with text embeddings before feeding them into the LLM decoder. A common approach is to project the audio feature information into the LLM’s text feature space through a connector [30, 26]. Specifically, the tensor containing audio features typically has its own distinct feature dimension. The audio feature embedding $\mathbf{Z}_{a}$ is projected into a vector with the same dimensionality as the text embedding using a projection function $w(\cdot)$: $\mathbf{H}_{a}=w(\mathbf{Z}_{a})$, where $\mathbf{H}_{a}$ denotes the projected audio embedding in the text embedding space. These audio embeddings are then concatenated with the input text embedding $\mathbf{H}_{t}$ to form a new embedding sequence that integrates both speech and text information: $\mathbf{H}=[\mathbf{H}_{a},\mathbf{H}_{t}]$. This combined sequence is then fed into the LLM. Some researchers have also implicitly incorporated the projection step within the original encoder, achieving modality projection by adjusting the encoder’s parameters during training [37]. Most models use a multi-layer perceptron (MLP) as the projection module, while some adopt more complex techniques such as Q-Former, which introduces trainable query tokens that attend to audio features and output fixed-length embeddings aligned with the LLM input space [26].
@@ IV-B2 Modality Fusion of Discrete Audio Tokens
    Token Mapping: The most intuitive approach is to map audio tokens to text tokens that the LLM can process. Tsunoo et al. [96] propose using CTC predictions to efficiently generate prompts—referred to as CTC prompts. After audio features are converted into text tokens using CTC, these tokens are merged with the tokenized input text to create a unified token sequence that represents both modalities. This sequence is then fed into the LLM for processing. This approach not only preserves the semantic content of the audio features but also ensures consistency in the LLM’s processing pipeline.
@@
IV-B3 Modality Alignment Strategies Discussion
Building on the preceding discussion, current Speech LLMs fuse modalities (i) under continuous audio representations via pre-decoder alignment or in-decoder cross-attention, and (ii) under discrete representations via token mapping or token-space expansion.
In practice, systems vary along several choices. Beyond fusion, a fundamental design choice is which audio representation to align. Using final-layer encodings maximizes semantic abstraction, while intermediate-layer features expose finer prosodic or phonetic cues. Attention-based pooling, learned query resamplers (e.g., Q-Former/Perceiver-style), strided subsampling, and CTC-based temporal compression are all used to control the rate of audio tokens fed to the LLM, balancing information retention and compute. It is worth noting that several recent studies have explored novel alignment paradigms. For example, LegoSLM [84] leverages LLM word embeddings to construct weighted representations of speech signals, while AlignFormer [99] effectively simplifies speech feature sequences by operating directly on the CTC posterior distributions. In a similar spirit, TASU [100] further exploits CTC posteriors to more effectively extract semantic information from audio and to balance the information rate between the speech and text modalities.
@@ IV-C LLM Inference
The LLM inference stage consists mainly of a text decoder, which is typically an autoregressive transformer decoder that generates text tokens conditioned on input audio and/or text. Architecturally, it often mirrors standard language models like GPT, but it is specially designed or adapted to handle multimodal inputs—such as projected audio embeddings or cross-attended audio features—alongside text tokens. To accommodate modality fusion, the decoder may incorporate special tokens (e.g., <aud>, <text>) or use prompt formatting to distinguish and contextualize different modalities during generation. Despite these adaptations, the decoder follows the same autoregressive objective: predicting the next text token given previous tokens and any additional input context. In models that generate both speech and text as output, the decoder may also be extended to support the generation of discrete audio tokens. One common approach, as seen in VALL-E, treats speech generation as a conditional codec language modeling task, where the model generates audio tokens conditioned on phoneme sequences and an acoustic prompt  [101, 102]. Another approach used in SpeechGPT expands the LLM’s vocabulary to include both text and speech tokens, enabling the model to generate audio token sequences directly alongside or following text, depending on the instruction [60]. In both cases, the generated audio tokens are passed through a neural codec decoder (e.g., EnCodec[58] or HiFi-GAN[103]) to reconstruct the final waveform.
@@ V Training Strategies
TABLE II: The following table shows the training strategy of Speech LLM. We show the training parameters at different stages, corresponding to the training data task type (as indicated inside the parentheses) and the type of capabilities the final model has. The – symbol indicates that the paper does not mention it in detail, while ✘ symbol indicates that the model does not have this stage.
@@ V-A Modality Alignment
At the outset, Speech LLMs must learn a shared representation between spoken input and text. In this modality alignment phase, models typically leverage vast speech data to bridge the audio-text gap. A common strategy is to incorporate a pre-trained speech encoder and align its acoustic representations with a text-based LLM. For example, Qwen-Audio begins with a pretrained Whisper encoder coupled to an LLM backbone via a projector [104]. The alignment is refined by supervised tasks that force audio and text to correspond. Notably, automatic speech recognition (ASR) and audio captioning (AAC) are usually used as fundamental alignment tasks [29, 27].
@@
The total parameters $\theta$ of the model are composed of $\theta_{\text{LLM}}$, $\theta_{\text{encoder}}$, and $\theta_{\text{adapter}}$ . The model is trained to map an input audio sequence $a$ to a textual output $y$ (transcript or caption). The training data consist of many triples $(a,p,y)$: an audio input $a$, a textual prompt or task descriptor $p$, and the expected textual output $y$. The model is optimized to predict $y$. In this stage, the textual prompt $p$ typically remains unchanged, which makes modality alignment easier to train [68, 105]. The training loss can be represented as
$\mathcal{L}_{\text{align}}(\theta)=-\log P(y\mid\theta_{\text{LLM}}(\theta_{\text{adapter}}(\theta_{\text{encoder}}(a)),p).$ (3)
@@
Depending on the training objectives, many models focus on a single speech-to-text task (e.g., speech recognition) rather than training a multitask speech model capable of fully understanding speech. Therefore, many models only have one stage of modality alignment. The most common approach is to train only the adapter [106, 105]. However, after the initial alignment, continuing to add LoRA parameters for further training has been shown to maintain better modality alignment and achieve superior alignment performance [68]. Another training approach involves simultaneously training the encoder, adapter, and LoRA parameters, as seen in SLAM and FireRedASR [107, 31]. Interestingly, both models employ Conformer as the encoder, and the FireRedASR model has achieved SOTA performance on current recognition leaderboards—perhaps demonstrating that unfreezing the encoder and training with more parameters can yield better alignment results. There is currently no unified consensus on the order of training parameters. A common practice is to first train the adapter, then train the encoder and adapter, and finally add LoRA parameters for joint training [33]. However, it is also advisable to include stages where some parameters are frozen during training [108]. For multitask models like Qwen-audio [37] and Seed-ASR [27], LoRA parameters of LLMs are typically not used as training parameters for modality alignment—partly to preserve the original powerful contextual and reasoning abilities of the LLMs, expecting them to play a role in subsequent multitask training. Not all models follow this pattern, however: SALMONN [26] trains LoRA parameters during the alignment phase.
@@ V-B Multitask Training
After modality alignment, speech features are mapped into the text space of the LLM. However, even when an LLM can accurately transcribe speech, it often loses the ability to reason about or follow instructions related to the transcribed content. Therefore, the LLM must undergo a multitask training phase to become a multitask speech model.
Taking models focused solely on ASR tasks as an example, during this phase many models add LoRA parameters to leverage the LLM’s contextual capabilities, including hotword usage and historical transcripts, to improve recognition results [109, 27]. However, these are not standard multitask training approaches but rather extensions of ASR tasks.
@@
The mainstream training approach resembles modality alignment: natural language instructions are concatenated with the aligned speech features and target text, and the model is directly trained for multitask via cross-entropy loss. Generally, LoRA parameters are added to the training at this stage to fine-tune the LLM, aiming to enable the LLM to acquire the multitask speech capability.
Notably, Qwen-Audio [37] employs a Whisper-like multitask training framework that uses task prefixes and does not use LoRA for fine-tuning. In contrast, OSUM [33] uses an ASR+X approach, where X represents other tasks. This means OSUM transcribes speech into text before performing any other audio tasks.
@@ V-C Preference Alignment
After multitask alignment, some advanced Speech LLMs apply preference alignment using reinforcement learning  [104]. This step optimizes the model’s responses according to human feedback or a learned reward function, refining qualities such as factual accuracy, relevance, or user satisfaction [110, 111], for example, Seed-ASR[27]. A prominent method is reinforcement learning from human feedback (RLHF), operationalized via Proximal Policy Optimization (PPO) [112]. In RLHF, the model is treated as a policy $\pi_{\theta}(y|X)$ that generates a response $y$ to an input $X$, and a reward model $R(X,y)$ (trained from human preference rankings) scores the response. The goal is to maximize expected reward $E_{X}[E_{y\sim\pi_{\theta}}[R(X,y)]]$ while keeping the language model’s output distribution close to the pre-trained model to avoid degrading fluency. PPO achieves this by updating $\theta$ in the direction of the policy gradient $\nabla_{\theta}J\approx E_{X,y}[R(X,y)\,\nabla_{\theta}\log\pi_{\theta}(y|X)]$ with constraints on the KL-divergence from the original policy. In practice, the PPO update uses a clipped surrogate objective to ensure stable improvements. An alternative, lightweight approach is Direct Preference Optimization (DPO) [113], which was employed in Qwen2-Audio’s final training. Unlike methods that require training a separate reward model, DPO directly leverages preference pairs $(y^{+},y^{-})$ for a given prompt $x$. It optimizes the policy to assign higher probability to the preferred response $y^{+}$ over the less favored $y^{-}$. The DPO loss can be formulated as:
@@
where $\sigma$ is the sigmoid function and $\beta$ is a temperature hyperparameter. By minimizing this loss, the model is encouraged to prefer $y^{+}$ over $y^{-}$, effectively aligning its outputs with human preferences without the complexity of a reward model training loop. By the end of this reinforcement or preference alignment phase, the Speech LLM is highly tuned to follow instructions in a human-preferred way: it will listen to a user’s voice request and respond with appropriate content. Crucially, this stage often improves the model’s factuality and safety by penalizing incorrect or undesirable continuations. Qwen2-Audio’s authors note that after DPO fine-tuning, the model more faithfully adheres to user intent and makes fewer mistakes in complex audio-centric instructions. The strength of the instruction alignment phase lies in making the model not just capable, but also reliably usable: it transforms the multitask-trained foundation into a polite conversational agent that can handle open-ended voice interactions. However, it is worth noting that this phase has received limited attention in current Speech LLM research focused on speech understanding. Most existing works primarily emphasize modality alignment and task-specific performance, while instruction-following behavior, dialogue consistency, and safety alignment remain underexplored. As Speech LLMs continue to evolve toward general-purpose audio agents, further investigation into instruction and preference alignment is expected to play a crucial role in improving usability, controllability, and user interaction quality.
@@ V-D Training of Speech LLMs with Discrete Sequence Modeling
In contrast to the three-stage training pipeline of continuous embedding models, Speech LLMs based on discrete speech tokens adopt a simpler and more unified training strategy. For these models (e.g., AudioPaLM [97], SpeechGPT [60], VALL-E [101]), after generating the discrete tokens for audio representations, these token sequences are treated analogously to text tokens, allowing a pretrained or newly initialized language model to be trained directly using a standard autoregressive objective.
Training typically proceeds by fine-tuning on paired speech-text data in a unified token space, without the need for a separate modality alignment stage, since discrete-token models do not introduce structural modifications such as adapters, which are common in continuous embedding models. This simplifies the pipeline and enables models to benefit directly from pretrained LLMs. In practice, alignment between modalities is learned implicitly through multitask fine-tuning, where the model learns to interleave and reason over both speech and text tokens. For example, SpeechGPT [60] introduces a “Paired Speech-Text Pretraining” phase to warm up the model on interleaved sequences, followed by instruction fine-tuning using synthetic multimodal dialogues. Unlike continuous-embedding Speech LLMs, mainstream discrete-token Speech LLMs typically skip explicit preference alignment (e.g., RLHF or DPO). Their training focuses on supervised multitask fine-tuning over speech–text data without a separate reward modeling stage. While emerging research such as SpeechAlign [114] has begun to explore preference optimization for codec-based language models, these methods remain experimental and are not yet part of the standard speech-LLM training toolbox.
@@ VI Datasests
In parallel with model advancements, the availability and design of datasets have also evolved to support increasingly complex speech understanding tasks. In this section, we categorize and analyze these datasets through the lens of their respective understanding objectives, as defined in II-B. We categorize speech understanding tasks according to their objectives, from simpler and more immediate tasks, such as perception tasks, to more complex and deeper objectives that require advanced reasoning, including shallow and deep cognition tasks. This progression parallels the evolution of speech models, which initially focused on handling fundamental tasks but have since advanced to address more intricate cognitive functions. Here, we provide a detailed overview of the various categories of speech datasets and examine how they have evolved, particularly in response to the development of LLMs and TTS.
@@ VI-A Datasets for Perception and Shallow Cognition Tasks
Perception and shallow cognition tasks focus on extracting information from audio signals and performing intuitive understanding, without requiring deep semantic reasoning. Tasks such as Automatic Speech Recognition (ASR), Speaker Diarization (SD), and Keyword Spotting (KWS) are examples of perception tasks, while emotion recognition and speech translation are considered shallow cognition tasks. These types of tasks correspond to traditional single-task datasets that predate the era of LLMs. A significant portion of the datasets used to train Speech LLMs consists of audio-text pairs, which are readily available from many traditional single-task datasets. In addition to these audio-text pairs, tasks like speaker diarization and keyword spotting often include additional metadata, such as timestamps and speaker labels, to enhance task performance. Rather than being limited to single-task datasets, multi-task datasets with rich metadata are also widely used, enabling training across multiple tasks. The AMI Meeting Corpus is a representative example of such datasets, which can be used for a variety of tasks as shown in [140, 141, 142, 143]. These datasets have become foundational in speech understanding due to their early development and wide availability. In addition to being used for pre-training Speech LLMs, these datasets can also serve as direct resources for downstream tasks, facilitating the fine-tuning of Speech LLMs.
Table III categorizes datasets associated with perception and shallow cognition tasks, including ASR, speaker analysis, KWS, emotion classification, and speech translation. This categorization highlights the diversity of task types, language coverage, and dataset scale, reflecting the fundamental role of these resources in the broader field of speech understanding.